Fermi National Laboratory
 

Accelerator Update for July 25, 2007 - July 27, 2007

 

Wednesday July 25

The day shift began with Operations monitoring store 5577 (50.76E30), the Recycler maintaining a stash (151.1E10), with the Antiproton Source (Pbar) stacking (42.0E10), with the Linac, Booster, and Main Injector (MI) operating normally, with MTest and MiniBooNE taking beam.

Operations terminated store 5577 at 3:02 PM.

Tevatron (Tev) experts began between store studies at 3:14 PM.

Operations reported at 4:21 PM that the Recycler had stashed antiprotons 3 times during the day shift.

MTest turned off at 6:11 PM

Operations put the Tevatron (TeV) into shot setup at 8:14 PM.

Operations established store 5580 at 9:57 PM with an initial luminosity of 228.3E30.

Operations reported at 11:58 PM that the Recycler had stashed antiprotons twice during the evening shift.

Thursday July 26

The midnight shift began with Operations monitoring store 5580 and the stash, with Pbar stacking, and with MiniBooNE taking beam.

MTest turned on at 7:18 AM.

Operations reported at 8:00 AM that the Recycler had stashed antiprotons four times during the midnight shift.

At 12:34 PM store 5580 was lost due to a Tevatron quench at D3. The quench was caused by a vacuum problem at D3.

CDF and D0 began controlled accesses into their halls at 1:30 PM.

Tev access began at 2:45 PM to repair a stuck kautzky valve at D3.

Operations reported at 3:50 PM that the Recycler had stashed antiprotons twice during the day shift.

Tev, CDF, and D0 accesses were completed by 4:17 PM. Quench recovery is continuing.

MTest turned off at 6:02 PM

Operations put the Tevatron (TeV) into shot setup at 10:16 PM.

Friday July 27

The midnight shift began with Operations in shot setup and MiniBooNE taking beam.

Operations established store 5582 at 12:50 AM with an initial luminosity of 203.35E30.

Linac RF station #3 tripped at 6:24 AM, experts were called and are investigating.

Operations reported at 7:59 AM that the Recycler had stashed antiprotons three times during the midnight shift.

The Plans for Friday and the weekend

The plan for today and for the weekend is to repair LRF3, continue to stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/27/2007    email Fermilab